OpenStudy (anonymous):

Could someone please help me get this one started? Find an equation of the line that passes through the points (1,4) and (-7,-4)

hero (hero):

1. Insert each point into y = mx + b, creating two equations.

OpenStudy (anonymous):

How do I figure which number to plug in where in the equation?

hero (hero):

Because (x,y) = (1,4) (x,y) = (-7,-4)

hero (hero):

So 4 = 1m + b -4 = -7m + b

hero (hero):

2. Isolate b in both equations

OpenStudy (anonymous):

ok, let me try this...lol

OpenStudy (anonymous):

ok. I'm drawing a blank... been a long day

hero (hero):

Did you isolate b in both equations yet?

OpenStudy (anonymous):

I'm having a brain fart as to how that is done and cannot find it in my textbook.

OpenStudy (anonymous):

http://www.purplemath.com/modules/strtlneq2.htm

hero (hero):

Isolate b in both equations: Subtract 1m from both sides: 4 - 1m = b Add 7m to both sides: -4 + 7m = b

hero (hero):

3. Set expressions b = b, then solve for m

OpenStudy (anonymous):

y=x+3?

OpenStudy (anonymous):

crap

hero (hero):

What steps did you use to figure that out?

OpenStudy (anonymous):

after you reminded me how to isolate B, +3 was the outcome for both, so I plugged in the rest of the information and failed horribly from there.

hero (hero):

So you set b = b to get 4 - 1m = -4 + 7m then solved for m, correct?

OpenStudy (anonymous):

not at all like that...

hero (hero):

Well that is what you are supposed to do. I can't imagine you doing anything other than that.

OpenStudy (anonymous):

I'm just the bad at math. lol, but I'm trying it again

OpenStudy (anonymous):

ok, the only thing that I am seeing here is to simplify down to 5m and 3m, but I know that neither of those will get me anywhere near the equation that I need.

hero (hero):

So from there, I'm assuming that you added 4 to both sides, then added 1m to both sides to get 4 + 4 = 1m + 7m

hero (hero):

right?

OpenStudy (anonymous):

I would lie if I said yes

hero (hero):

Well, that's what you do. Can you solve for m from there?

OpenStudy (anonymous):

hang on, I'll show you how I did it

hero (hero):

Why would I want to see how you did it the wrong way?

OpenStudy (anonymous):

4-1m=-4+7m -4 +4 -1m+7m 6m then somehow on my work I had 5m+3m=8m

OpenStudy (anonymous):

I am awful at this stuff

hero (hero):

You don't add 4 on one side then subtract 4 on the other.

hero (hero):

What ever operation you choose to perform, you do that to BOTH sides.

OpenStudy (anonymous):

ok, so I don't clear to the left of the = in each one and then combine the numbers next to m?

hero (hero):

No, you use inverse operations so that you have like terms on each side.

hero (hero):

So you should have ended up with 4 + 4 on one side and 1m + 7m on the other side.

OpenStudy (anonymous):

so that gives me 8=8m^2

hero (hero):

You mean 8 = 8m You would only get m^2 if you multiplied , which you did not.

hero (hero):

If you see a plus sign between two expressions, you add, not multiply.

OpenStudy (anonymous):

That's right, forgot that rule.

hero (hero):

You forgot how to add?

OpenStudy (anonymous):

lol no, the rule of only getting m^2 if multiplication were used.

hero (hero):

Did you solve for m yet?

OpenStudy (anonymous):

not yet

hero (hero):

I can't imagine what is taking you so long.

OpenStudy (anonymous):

y=7x+7

hero (hero):

Okay, now you're just going off on a tangent

hero (hero):

This is what happens when you try to walk someone through a problem.

hero (hero):

Now, I'm just going to post the solution.

hero (hero):

1. Insert (1,4) and (-7,-4) into y = mx + b, creating two equations: 4 = 1m + b -4 = -7m + b 2. Isolate b in both equations to get: 4 - 1m = b -4 + 7m = b 3. Set expressions b = b to get: 4 - 1m = -4 + 7m 4. Add 4 to both sides, then add 1m to both sides: 4 + 4 = 1m + 7m 5. Combine like terms to get: 8 = 8m 6. Divide both sides by 8 to isolate and solve for m: 8/8 = m 1 = m 7. Insert the value of m into one of the equations in step 2: 4 - 1m = b 4 - 1(1) = b 4 - 1 = b 3 = b 8. The equation of the line is y = x + 3

hero (hero):

You got y = x + 3 the first time, but you couldn't show how you got it.

OpenStudy (anonymous):

I said that I was awful at this subject. I tried my best, and I am sorry if you lost patience with me. There is no need to be crass or insulting.

hero (hero):

I wasn't trying to be. I was just saying that I asked you to post your steps for how you got y = x + 3 the first time.